OpenStudy (undeaddragon26):

Name FIVE x-values for which the cosine function equals 0.

OpenStudy (johnweldon1993):

Hint: \(\large cos(x) = 0\)  as you can see the cosine function oscillates so it can be 0 at many points I'll give you one.....At x = 90 degrees or \(\large \frac{\pi}{2}\)

OpenStudy (undeaddragon26):

would it be 90 times every odd number

OpenStudy (johnweldon1993):

That will also play the same in the negative region since cos is an odd function we also know that \[\large cos(-\frac{\pi}{2}) = 0\]

OpenStudy (undeaddragon26):

90*1 90*3 etc ?

OpenStudy (johnweldon1993):

You are correct. 90 degrees times every odd number (positive or negative) will keep resulting of cos(x) = 0

OpenStudy (undeaddragon26):

ok thx